Given f(x) = 2x + 3, find the value of x where f(x) = 7. Do not include x = in your answer. Provide your answer below:
Answer #1

Geven have of a f(n) = 2n + 3 we to find tue value where f (n) =7. Now f (x)=7 mo, 22+3 = 7 2a= 7-3 = 4 2n=4 or, naa = 2 q An

The solution is given in details